# The incompatibility of the Condorcet winner and loser criteria with positive involvement and resolvability

### v5

@@ 3 Result
We prove the $n=1$ case. Reasoning as in the proof of Theorem 6 handles $n>1$. For positive involvement, consider a $\mathbf{Q}_{1}$ as in Figure 3, wherein $\{a,b,c,e\}$ is a set of clones. By Lemma 3, $F(\mathbf{Q}_{1})\subseteq\{a,d\}$. By the Condorcet winner criterion, $F(\mathbf{Q}_{1-b})=\{a\}$, so by independence of clones (part (i) or (ii)) and $F(\mathbf{Q}_{1})\subseteq\{a,d\}$, we have $a\in F(\mathbf{Q}_{1})$. Then obtain $\mathbf{Q}_{2},\dots,\mathbf{Q}_{5}$ from $\mathbf{Q}_{1}$ by adding and subtracting the profiles in Figure 4. The rest of the proof is just like that of Theorem 4 but with $\mathbf{Q}_{i}$ instead of $\mathbf{P}_{i}$. For negative involvement, by block preservation, we can add as many blocks of all linear orders as we wish to $\mathbf{Q}_{1}$ to obtain a $\mathbf{Q}_{1}^{+}$ with $F(\mathbf{Q}_{1})\subseteq F(\mathbf{Q}_{1}^{+})\subseteq\{a,d\}$ by the variant of Lemma 3 in the proof of Theorem 7. Hence $F(\mathbf{Q}_{1})\subseteq\{a,d\}$, which implies $a\in F(\mathbf{Q}_{1})$ as above and therefore $a\in F(\mathbf{Q}_{1}^{+})$. Then starting from $\mathbf{Q}_{1}^{+}$, we reason with negative involvement instead of positive involvement as in the proof of Theorem 7.∎
@@ Acknowledgements
I thank Diego Bejarano and Emma Boniface for helpful discussions of the problem posed at the end of Holliday 2024 and Dominik Peters for helpful feedback on this paper.

### v6

@@ 3 Result
We prove the $n=1$ case. Reasoning as in the proof of Theorem 6 handles $n>1$.
For part 1, consider a $\mathbf{Q}_{1}$ as in Figure 3, wherein $\{a,b,c,e\}$ is a set of clones. By Lemma 3, $F(\mathbf{Q}_{1})\subseteq\{a,d\}$. By the Condorcet winner criterion, $F(\mathbf{Q}_{1-b})=\{a\}$, so by independence of clones (part (i) or (ii)) and $F(\mathbf{Q}_{1})\subseteq\{a,d\}$, we have $a\in F(\mathbf{Q}_{1})$. Then obtain $\mathbf{Q}_{2},\dots,\mathbf{Q}_{5}$ from $\mathbf{Q}_{1}$ by adding and subtracting the profiles in Figure 4. The rest of the proof is just like that of Theorem 4 but with $\mathbf{Q}_{i}$ instead of $\mathbf{P}_{i}$.
For part 2, by block preservation we can add as many blocks of all linear orders as we wish to $\mathbf{Q}_{1}$ to obtain a $\mathbf{Q}_{1}^{+}$ with $F(\mathbf{Q}_{1})\subseteq F(\mathbf{Q}_{1}^{+})\subseteq\{a,d\}$ by the variant of Lemma 3 in the proof of Theorem 7. Hence $F(\mathbf{Q}_{1})\subseteq\{a,d\}$, which implies $a\in F(\mathbf{Q}_{1})$ as above and therefore $a\in F(\mathbf{Q}_{1}^{+})$. Then starting from $\mathbf{Q}_{1}^{+}$, we reason with negative involvement instead of positive involvement as in the proof of Theorem 7.∎
Thus, the tradeoff noted in Section 1 is shown by several inconsistent sets of axioms, summarized in Table 1.
@@
Condorcet Condorcet positive negative resolvability independence block
winner loser involvement involvement of clones preservation
Theorem 4 $circleddash \circleddash \circleddash$
Theorem 7 $circleddash \circleddash \circleddash$
Theorem 8.1 $circleddash \circleddash \circleddash$
Theorem 8.2 $circleddash \circleddash \circleddash \circleddash$
Table 1: summary of main impossibility theorems. For each theorem, the axioms that are jointly inconsistent are indicated by circled marks in that row. In each case, resolvability can be replaced by $n$-voter resolvability.
@@ Acknowledgements
I thank Diego Bejarano and Emma Boniface for helpful discussions of the problem posed at the end of Holliday 2024, as well as Dominik Peters and an anonymous referee for helpful feedback on this paper.